# Robust certification of quantum instruments through a sequential communication game

### v1

@@ Abstract
We propose a communication game in the sequential measurement scenario, involving a sender and two receivers with restricted collaboration among the latter parties. In the framework of the prepare-transform-measure scenario, we find a prominent quantum advantage in the receiver’s decoding of the message originally encoded by the sender. We show that an optimal trade-off between the success probabilities of the two receivers enables self-testing of the sender’s state preparation, the first receiver’s instruments, and the measurement device of the second receiver. Our protocol enables a more robust certification of the unsharp measurement parameter of the first receiver compared to the protocol without collaboration among the receivers. We further generalize our game to higher-dimensional systems, revealing greater quantum advantage with an increase in dimensions.
@@ I Introduction
The study of the performance of quantum systems for communication and computational purposes gave birth to quantum information science [1, 2]. The study was essential as the technology was maturing towards manipulating single quantum systems [3]. In present times, the certification of quantum systems is regarded as an important cornerstone for the successful and secure implementation of quantum technology protocols. Robust self-testing forms the basis of certification of quantum states and measurements in a device-independent manner [4, 5, 6, 7, 8, 9, 10]. Practically relevant extensions in semi-device scenarios have also been proposed [11, 12, 13, 14, 15, 16]. The certification of quantum instruments plays a key role in the prepare-transform-measure scenario, forming an essential component of quantum communication in sequential networks.
@@
On the other hand, the idea of employing sequential measurements in investigating foundational and practical information processing issues have gained much importance in recent years [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. The sequential Bell test scenario has been employed to harness unbounded randomness from a single pair of entangled states [57] and self-testing quantum instruments as well [58]. Moreover, sequential measurements have been employed to study various other quantum correlations such as quantum steering [59, 60, 61], and sequential communication protocols have been suggested for information processing tasks such as entanglement detection, teleportation, and remote state preparation [62, 63, 64].
@@
The generalization of the standard $2\rightarrow 1$ QRAC protocol to the sequential scenario has been performed [65, 66, 67], enabling various certification tasks. In [65], $2\rightarrow 1$ QRAC was extended to three parties where two independent receivers try to decode the bit position encoded earlier by the sender, randomly and sequentially. Here, the first decoder must have to measure weakly so that the second decoder can have a non-classical success probability. In this prepare-transform-measure (PTM) scenario, the success probability of the two decoders exhibits a trade-off, absent in the analogous classical counterpart. Such a protocol can be employed to self-test quantum instruments [65], and non-projective [68] and unsharp measurements [66] in a semi-device-independent manner. It may be noted that experimental implementations of sequential measurement protocols have also been performed [69, 70].
@@
In the present work, we propose a new communication game enabling a more robust certification of quantum instruments. The standard $2\rightarrow 1$ QRAC scenario is extended to a tripartite communication game where there are two sequential receivers who are not completely independent. Here, the sender (Aparna) and the first receiver (Barun) act in the same way as in the standard RAC task, whereas the second receiver (Chhanda) has to guess the remaining bit position so that at the end of the game, both receivers together may have the full information about the message encoded by the sender. Specifically, in the context of $2\rightarrow 1$ QRAC, two bits of information are encoded into one bit by Aparna. Barun decodes the bit value of some position according to his random input. In our protocol, we allow restricted collaboration between two decoding receivers to make the task non-trivial and interesting as well. After decoding the asked bit value, Barun passes only his input information and the post-measured system to Chhanda without leaking any output information. We compute the optimal classical success probability of this proposed communication game and demonstrate an advantage through a quantum strategy.
@@
We explore a trade-off between the success probability of Barun and Chhanda, which is absent in the classical case. The optimal trade-off leads to the certification of Barun’s quantum instrument together with Aparna’s state preparation and Chhanda’s measurement. We next investigate the robustness of the certification of quantum instruments in the presence of noise and estimate the lower and upper bounds on the sharpness parameter of Barun’s measurement. A significant outcome of our analysis is that our certification task is shown to be more robust compared to the approach of employing sequential QRAC without collaboration [65]. Since the gap between the lower and upper bounds of the estimated sharpness parameter of the first receiver’s measurement is narrower in our case, we are able to improve upon the earlier results in the context of our new communication game. We further generalize our game to the higher-dimensional $2^{d-1}\rightarrow 1$ QRAC scenario. We find the optimal classical success probability and numerically compute the quantum advantage up to dimension $6$.
@@
II Sequential game with restricted collaboration
Figure 1: Schematic diagram for a two-receiver communication game with restricted collaboration in ‘d’-dimensional quantum settings. Sender Aparna compressed a 2-dit message into a qudit and sends it to the first receiver, Barun, who decodes a dit position randomly. The task of the second receiver, Chhanda, is to guess the leftover information under the restriction that she can only use the knowledge of Barun’s input and has no information of his output from the communication.
In this section, we describe our proposed new communication game pertaining to the prepare-transform and measure scenario among three parties: Aparna, Barun, and Chhanda. Initially, two dits (higher-dimensional generalizations of a bit) of information are encoded into one dit by the sender, Aparna. The first receiver, Barun, tries to decode some partial message randomly from the encoded system and then passes it to Chhanda, who tries to retrieve the remaining message. The crucial point is that, unlike the protocol [65], here two receivers are not independent, and we allow some limited collaboration between them. In the next two subsections, we will discuss both the classical and quantum strategies to play this game, followed by results in the subsequent sections.
@@ II.1 Two-receiver classical access code
In the classical version of this game, Aparna first selects a message randomly from $d^{2}$ options, which is a 2-dit string $x=(x_{0},x_{1})$. She encodes it into a one-dit message ‘m’ and sends it to Barun where, $x_{i},~{}m\in\{0,1,...,d-1\}$. Barun receives a random input $y\in\{0,1\}$ and needs to guess the $x_{y}$ dit value of Aparna’s message. After decoding the asked dit value, Barun will let Chhanda know about which position she has to decode and relay the post-measured system with a restriction. Here, the restriction is that the relayed system will carry no information about Barun’s output, i.e., Barun passes only his input information and the post-measured system to Chhanda without leaking any output information. Chhanda’s task is to infer the other dit, $x_{\bar{y}}$, of Aparna’s message (see, Fig.1).
In this communication task, Barun’s role is similar to that of the traditional RAC [21, 22]. To maximize Barun’s success probability of correctly retrieving the asked dit value, they can adopt the first-dit encoding strategy without loss of generality. With this approach, when Barun’s input is $y=0$, he retrieves the dit with probability 1. However, for the second dit (when $y=1$), his success probability is $\frac{1}{d}$ as he has no option other than a random guess. Therefore, on average, his success probability is $\frac{1}{2}\left(1+\frac{1}{d}\right)$. This classical strategy was conjectured to be optimal [22] and proved recently in [71].
Chhanda, however, lacks this advantage due to the restricted communication from Barun. As she only has the information about Barun’s input but not about his output, her success probability remains limited to $\frac{1}{d}$. The total classical success probability of retrieving the sender’s full encoded message, jointly by the two receivers, according to the above stated strategy of restricted collaboration, is given by,
@@ II.2 Two-receiver quantum access code
Now let’s describe the two receiver quantum access code with limited collaboration between the receivers. Aparna prepares a quantum state $\rho_{x}$ that encodes the selected classical message, $x=(x_{0},x_{1})$$\in\{0,1,...d-1\}^{2}$. Two receivers, Barun and Chhanda, only know the set of states of dimension $d$ from which Aparna selects in a single run of the game, but not the exact state. After preparing the state, Aparna sends it to Barun, who applies his measuring instrument, represented mathematically by the Kraus operators $\{K_{b|y}\}$. When this instrument acts on the state $\rho_{x}$, the outcome is $b\in\{0,1,\dots,d-1\}$, and the updated state according to the Lüdder transformation rule [72], is given by,
@@
To satisfy the completeness relation for Barun’s measurement, we require that $\sum_{b=0}^{d-1}M_{b|y}=\mathbb{I}$, where $M_{b|y}=K_{b|y}^{\dagger}K_{b|y}$ are the elements of a positive operator-valued measure (POVM). The correlation between Aparna and Barun is captured by the average success probability of correctly guessing the asked dit value, which can be obtained via Born’s rule in the following way. In the present work, we use the words average success probability, success probability, and figure of merit of the game interchangeably.
@@
According to the rules of the game, Barun will disclose his measurement choice, denoted by $y$, but not the corresponding outcome. Therefore, the post-measurement state received by Chhanda, is simply the average over all possible outcomes for a fixed input $y$, which is
@@
Now, to retrieve the leftover information, Chhanda will perform a generalized measurement $\{N_{0},N_{1}\}$ on $\tilde{\rho}_{x}^{y}$, where, $\sum_{b=0}^{d-1}N_{b|y}=\mathbb{I}$. Chhanda’s figure of merit is the success probability of correctly guessing the remaining part of the encoding and is given by the following way.
@@
The total average success probability of retrieving the full information of Aparna’s encoding by Barun and Chhanda together employing this restricted collaborative strategy is
@@ III Quantum advantage and robust certification of instruments
In this section, we analyze the one sender, two receiver communication game in the presence of restricted collaboration between the receivers. We show explicitly that in quantum settings, an interesting non-trivial trade-off between the figure of merit of the two receivers arises, which leads to robust certification of quantum instruments. We consider the minimal scenario here, i.e., a two-bit message compressed into one bit, and generalize it to higher dimensions in the next section.
@@ III.1 Trade-off between Barun’s and Chhanda’s respective figure of merit
$\forall y,~{}~{}\sum_{x}\tilde{\rho}_{x}^{y}=\frac{\mathbb{I}}{2},$
and $\text{and }P_{AB}^{Q}=\beta.$
@@
In the above formulated SDP (7), we maximize Chhanda’s success probability by optimizing over all possible state preparations and measurements. The condition $P_{AB}^{Q}=\beta$ is enforced using suitable instruments, and its Kraus operator representation in the polar decomposition form is given as $K_{b|y}=U_{y|b}\sqrt{M_{b|y}}$, where $U_{y|b}$ represents a unitary operator, and $\{M_{b|y}\}$ is a POVM. The second last constraint is to impose the restriction that Chhanda has only the information about Barun’s input but no way to know his output.
@@
[…]
@@ III.2 Self-testing based on optimal quantum strategy
[…]
@@ III.3 Robust certification of sharpness parameter
In realistic conditions, all quantum systems deviate from the ideal target description, and hence, less than optimal success probabilities are observed, which can’t pinpoint the desired objects. Robustness analysis is the estimation of those objects for the suboptimal observed values. Therefore, any meaningful self-testing argument must come with a robustness analysis. Our above-mentioned self-testing statements in Sec.III.2 are valid only when the optimal trade-off point is achieved, which is not the case in all practical experimental conditions. Here we specifically consider the sharpness parameter of Barun’s measurement, which characterizes his instrument. We show how one can certify the sharpness parameter $\eta$, of Barun’s measurement in the presence of noise, in a more robust way than that in the approach based on sequential QRAC without collaboration [65]. We provide upper and lower bounds on the sharpness parameter for less than optimal observed success probabilities, which merge to the same value in noiseless conditions. Let’s state the result briefly here, and for the detailed analysis, we refer to the Appendix (B).
@@
[…]
$\eta\geq\sqrt{2}(2P_{AB}^{Q}-1).$ (10)
This bound is non-trivial whenever $P_{AB}^{Q}>1/2$. On the other hand, from the requirement that $P_{AC}^{Q}$ is non-classical, we can place an upper bound on $\eta$, given by
$\eta\leq\sqrt{8P_{AC}^{Q}-8(P_{AC}^{Q})^{2}-1}.$ (11)
@@
From the observed value of $(P_{AB}^{Q},P_{AC}^{Q})$, the pair, one can estimate the sharpness parameter, which lies within the upper and lower bounds. The narrower the gap between the two bounds, the better the certification. Specifically, we consider a noise model where Aparna’s preparation is given by $\rho_{x}^{{}^{\prime}}=p_{1}\rho_{x}+(1-p_{1})\frac{\mathbb{I}}{2}$, and the POVM elements corresponding to the measurement devices of Barun and Chhanda are represented as $M_{y}^{{}^{\prime}}=p_{2}M_{y}+(1-p_{2})\frac{\mathbb{I}}{2}$ and $N_{z}^{{}^{\prime}}=p_{3}N_{z}+(1-p_{3})\frac{\mathbb{I}}{2}$, respectively. Here, $p_{1}$ denotes the visibility of Aparna’s preparation device, $p_{2}$ and $p_{3}$ the visibilities of Barun’s and Chhanda’s measurement devices, respectively. In Fig. 3, we graphically illustrate the width of the plausible range for the sharpness parameter associated with the underlying noisy measurements.
In a noisy and suboptimal scenario, the sharpness parameter can be estimated within a significantly narrower range using our proposed communication game compared to the estimation based on the sequential QRAC protocol with no collaboration among the receivers [65]. For instance, suppose Aparna’s preparations have a visibility of 95%, while Barun’s and Chhanda’s measurements achieve visibilities of 93% and 95%, respectively. Under these conditions, our protocol estimates the sharpness parameter with an error gap of approximately $0.1133$, whereas the method in Ref. [65] leads to a larger error gap of $0.1964$. In Fig. 3(b), for the same visibility values, we plot the gap $(\delta,\delta^{*})$ as a function of $\eta^{\text{Target}}$, where $\eta^{\text{Target}}$ denotes the value of the sharpness parameter that an experimentalist may want to self-test. Detailed calculations and comparison plots for two other noisy scenarios are provided in Fig. 5 of Appendix B.
In Table 1, we present a comparison of the gap between the upper and lower bounds, estimated from the observed figures of merit in our communication game with restricted collaboration between the receivers and the protocol without such collaboration [65], for three different sets of noise levels. The gap between the bounds over the entire range of allowed observed success probabilities for the visibility set $(p_{1}=0.95,p_{2}=0.9,p_{3}=0.95)$ for the first set of noise levels is illustrated in Fig. 3(a). The results for the other two noisy scenarios are extracted from Fig. 5 in Appendix B. It can be seen that the protocol for self-testing the sharpness parameter based on restricted collaboration proposed by us is more robust compared to the one in Ref. [65] for a range of parameter values.

### v2

@@ Abstract
We propose a communication game in the sequential measurement scenario, involving a sender and two receivers with restricted communication among the latter parties. In the framework of the prepare-transform-measure scenario, we find a prominent quantum advantage in the receiver’s decoding of the message originally encoded by the sender. We show that an optimal trade-off between the success probabilities of the two receivers enables self-testing of the sender’s state preparation, the first receiver’s instruments, and the measurement device of the second receiver in a semi device-independent way. Our protocol enables a more robust certification of the unsharp measurement parameter of the first receiver compared to an earlier protocol. We further generalize our game to higher-dimensional systems, revealing greater quantum advantage with an increase in dimensions.
@@ I Introduction
The study of the performance of quantum systems for communication and computational purposes gave birth to quantum information science [1, 2]. The study was essential as the technology was maturing towards manipulating single quantum systems [3]. In present times, the certification of quantum systems, is regarded as an important cornerstone for the successful and secure implementation of quantum technology protocols. In the literature, various types of certification schemes are classified according to the set of assumptions and level of trust required for the experimental implementation. The certification scheme is called robust if it remains meaningful in presence of some amount of deviation from the ideal description. Robust self-testing usually refers to the certification of quantum states and measurements in a device-independent manner [4, 5, 6, 7, 8, 9, 10]. In semi-device independent scenarios, requirements are lesser stringent, which is practically more relevant [11, 12, 13, 14, 15, 16]. Witnesses which require trust on preparation and measurement devices, belong to the category of device-dependent formalism [17, 18].
@@
On the other hand, the idea of employing sequential measurements in investigating foundational and practical information processing issues have gained much importance in recent years [48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. The sequential Bell test scenario has been employed to harness unbounded randomness from a single pair of entangled states [59] and self-testing quantum instruments as well [60]. Moreover, sequential measurements have been employed to study various other quantum correlations such as quantum steering [61, 62, 63], and sequential communication protocols have been suggested for information processing tasks such as entanglement detection, teleportation, and remote state preparation [64, 65, 66]. Here we refer to the non-sequential setting as the standard one where certification of quantum instruments, non-projective and unsharp measurements is not possible. This can be circumvented invoking sequential settings [67, 68, 69]. The certification of quantum instruments, which plays a key role in quantum science and technology, is the main focus of the present paper.
@@
The generalization of the standard $2\rightarrow 1$ QRAC protocol to the sequential scenario has been performed [67, 68, 70], enabling various certification tasks. In [67], $2\rightarrow 1$ QRAC was extended to three parties where two independent receivers try to decode the bit position encoded earlier by the sender, randomly and sequentially. Here, the first decoder must have to measure weakly so that the second decoder can have a non-classical success probability. In this prepare-transform-measure (PTM) scenario, the success probability of the two decoders exhibits a trade-off, absent in the analogous classical counterpart. Such a protocol can be employed to self-test quantum instruments [67], and non-projective [69] and unsharp measurements [68] in a semi-device-independent manner going beyond the standard settings of certification of quantum states and measurements. It may be noted that experimental implementations of sequential measurement protocols have also been performed [71, 72].
@@
In the present work, we propose a new type of communication game and show its quantum advantage, which enables certification of various quantum components in a semi device-independent way. Surprisingly, our scheme yields a more robust certification of quantum instruments, known so far. The standard $2\rightarrow 1$ QRAC scenario is extended to a tripartite communication game where there are two sequential receivers who are not completely independent, unlike the previous extension [67]. Here, the sender (Aparna) and the first receiver (Barun) act in the same way as in the standard RAC task, whereas the second receiver (Chhanda) has to guess the remaining bit position so that at the end of the game, both receivers together may have the full information about the message encoded by the sender. Specifically, in the context of $2\rightarrow 1$ QRAC, two bits of information are encoded into one bit by Aparna. Barun decodes the bit value of some position according to his random input. After decoding the asked bit value, Barun passes the post-measured system to Chhanda without leaking any output information, and the task of Chhanda is to guess the other bit value. We compute the optimal classical success probability of this communication game and demonstrate a prominent advantage through a quantum strategy.
@@
Next, we explore a trade-off between the success probability of Barun and Chhanda, which is absent in the classical case. The optimal trade-off leads to the certification of Barun’s quantum instrument together with Aparna’s state preparation and Chhanda’s measurement. We then investigate the robustness of the certification of quantum instruments in the presence of noise and estimate the lower and upper bounds on the sharpness parameter of Barun’s measurement. A significant outcome of our analysis is that our certification task is shown to be more robust compared to the approach of employing sequential QRAC without collaboration [67]. Since the gap between the lower and upper bounds of the estimated sharpness parameter of the first receiver’s measurement is narrower in our case, we are able to improve upon the earlier results in the context of our new communication game. We further generalize our game to the higher-dimensional $2^{d-1}\rightarrow 1$ QRAC scenario. We find the optimal classical success probability and numerically compute the quantum advantage up to dimension six.
@@
II Two receiver sequential game with restricted communication
Figure 1: Schematic diagram for a two-receiver communication game with restricted collaboration in ‘d’-dimensional quantum settings. Sender Aparna compressed a 2-dit message into a qudit and sends it to the first receiver, Barun, who decodes a dit position randomly. The task of the second receiver, Chhanda, is to guess the leftover information under the restriction that she does not have Barun’s output information except wild guess.
In this section, we describe our proposed new communication game pertaining to the prepare-transform and measure scenario among three parties: Aparna, Barun, and Chhanda. Initially, two dits (higher-dimensional generalizations of a bit) of information are encoded into one dit by the sender, Aparna. The first receiver, Barun, tries to decode some partial message according to the random input, obtained from some referee, by measuring the encoded system and then passes it to Chhanda, who tries to retrieve the remaining message. The crucial point is that, unlike the protocol [67], here two receivers are not independent, and Barun’s communication is such that Chhanda can not have his output information except wild guess. In the next two subsections, we will discuss both the classical and quantum strategies and relevant pay-off functions of this game, followed by results in the subsequent sections.
@@ II.1 Two-receiver classical access code
In the classical version of this game, Aparna first selects a message randomly from $d^{2}$ options, which is a 2-dit string $x=(x_{0},x_{1})$. She encodes her two dits $(x_{0},x_{1})\in\{0,1,\dots,d-1\}^{2}$ into a one-dit message $m\in\{0,1,\dots,d-1\}$ and sends it to Barun. Barun and Chhanda receive a common random input $y\in\{0,1\}$, say, from a referee. Barun’s task is to guess the dit $x_{y}$, while Chhanda’s task is to guess the complementary dit $x_{\bar{y}}$. After producing his guess, Barun forwards the post-measurement system to Chhanda, satisfying the constraint that the system carries no information about his output. Then Chhanda attempts to infer $x_{\bar{y}}$ (see Fig. 1) using the system.
The performance of this communication task is evaluated via three payoff functions. They are quantified by the average success probability of correctly guessing the asked bit value individually by Barun, Chhanda and that of by them jointly.
First payoff: The average success probability of Barun is given by,
$P_{AB}^{C}(b=x_{y}|x,y)$ $=\frac{1}{2d^{2}}\sum_{x,y}p(b=x_{y}|x,y)$ (1)
In this task, Barun’s role is similar to that of the traditional RAC [23, 24]. To maximize Barun’s success probability of correctly retrieving the asked dit value, they can adopt the first-dit encoding strategy without loss of generality. With this approach, when Barun’s input is $y=0$, he retrieves the dit with probability 1. However, for the second dit (when $y=1$), his success probability is $\frac{1}{d}$ as he has no option other than a random guess. Therefore, on average, his success probability ($P_{AB}^{C}$) is $\frac{1}{2}\left(1+\frac{1}{d}\right)$. This classical strategy was conjectured to be optimal [24] and proved recently in [73].
Second payoff: Similarly, the average success probability of Chhanda is given by,
$P_{AC}^{C}(c=x_{\bar{y}}|x,y)$ $=\frac{1}{2d^{2}}\sum_{x,\bar{y}}p(c=x_{\bar{y}}|x,\bar{y}).$ (2)
Unlike Barun, Chhanda can’t do better than a wild guess. Since she receives only the post-measurement system from Barun which carries no information about his output. Thus, her success probability $(P_{AC}^{C})$ is limited to $\tfrac{1}{d}$.
Third payoff: The total classical success probability of retrieving the sender’s full encoded message, jointly by the two receivers, according to the above stated strategy of restricted collaboration, is given by,
@@ II.2 Two-receiver quantum access code
Now let’s describe the two receiver quantum access code i.e., quantum version of the proposed game. Aparna prepares a quantum state $\rho_{x}$ that encodes the selected classical message, $x=(x_{0},x_{1})$ $\in\{0,1,...d-1\}^{2}$. Two receivers, Barun and Chhanda, only know the set of states of dimension $d$ from which Aparna selects in a single run of the game, but not the exact state. After preparing the state, Aparna sends it to Barun, who applies his measuring instrument, represented mathematically by the Kraus operators $\{K_{b|y}\}$. When this instrument acts on the state $\rho_{x}$, the outcome is $b\in\{0,1,\dots,d-1\}$, and the updated state according to the Lüdder transformation rule [74], is given by,
@@
To satisfy the completeness relation for Barun’s measurement, we require that $\sum_{b=0}^{d-1}M_{b|y}=\mathbb{I}$, where $M_{b|y}=K_{b|y}^{\dagger}K_{b|y}$ are the elements of a positive operator-valued measure (POVM). The correlation between Aparna and Barun is captured by the average success probability of correctly guessing the asked dit value, which constitute the first payoff and is given by,
@@
In the present work, we use the words average success probability, success probability, and figure of merit of the game interchangeably. According to the rules of the game, Barun only passes his post-measured state without leaking output information. Therefore, the post-measurement state received by Chhanda, is given by the average over all possible outcomes for a fixed input $y$, which is,
@@
Now, to retrieve the leftover information, Chhanda will perform a generalized measurement $\{N_{0},N_{1}\}$ on $\tilde{\rho}_{x}^{y}$, where, $\sum_{b=0}^{d-1}N_{b|y}=\mathbb{I}$. Chhanda’s figure of merit is the success probability of correctly guessing the remaining part of the encoding, which is the second payoff function of the game, and computed as given bellow.
@@
Our third payoff, the total average success probability of retrieving the full information of Aparna’s encoding by Barun and Chhanda together employing this restricted collaborative strategy is
@@
If $P_{AB}^{Q}>P_{AB}^{C}$ and $P_{AC}^{Q}>P_{AC}^{C}$, then both Barun and Chhanda obtain a quantum advantage and similar conclusion holds for joint success probability when $P^{Q}>P^{C}$ is obtained.
@@ III Quantum advantage and robust certification of instruments
In this section, we analyze quantum strategy of the one-sender, two receiver new communication game. We show explicitly that in quantum settings, an interesting non-trivial trade-off between the first and the second payoff function, which leads to robust certification of quantum instruments along with SDI certification of states and measurements. We consider the minimal scenario here, i.e., a two-bit message compressed into one bit, and generalize it to higher dimensions in the next section.
@@ III.1 Trade-off between Barun’s and Chhanda’s respective figure of merit
$\forall y,~~\frac{1}{4}\sum_{x}\tilde{\rho}_{x}^{y}=\frac{\mathbb{I}}{2},$
$\text{and }P_{AB}^{Q}=\beta.$
@@
In the above SDP (9), we maximize Chhanda’s quantum success probability by optimizing over all possible state preparations and measurement strategies. The first constraint ensures that Aparna’s encodings $\rho_{x}$ are valid qubit states. The second and third constraints specify that both Barun’s and Chhanda’s measurements are valid POVMs, with Barun’s instruments further expressed in Kraus form $K_{b|y}=U_{y|b}\sqrt{M_{b|y}}$, where $U_{y|b}$ is a unitary and $\{M_{b|y}\}$ a POVM. The condition $P_{AB}^{Q}=\beta$ is imposed through these instruments to fix Barun’s performance level. The second last constraint $\frac{1}{4}\sum_{x}\tilde{\rho}_{x}^{y}=\tfrac{\mathbb{I}}{2}$ enforces the restriction that the system relayed by Barun contains no information about his output, ensuring that Chhanda only has access to Barun’s input but not his outcome.
@@
The trade-off between $P_{AB}^{Q}$ and $P_{AC}^{Q}$ is displayed in the Fig.2. The above quantum strategy is actually optimal, since it saturates the upper bound on $P_{AC}^{Q}$ for any chosen value of $P_{AB}^{Q}=\beta$ (see Appendix (A)). Next, we move on to show explicitly how the optimal success probability point on the quantum boundary of the trade-off curve in correlation space leads to the certification of various quantum components involved in the strategy in a semi-device-independent way.
@@ III.2 Self-testing based on optimal quantum strategy
For the purpose of self-testing, we need to identify the underlying quantum objects uniquely up to some symmetry operations. The optimal trade-off relation allows us to certify Aparna’s state preparation, Barun’s instruments, and Chhanda’s measurements within a collectively chosen reference frame. For more detailed proof of the optimal trade-off and self-testing argument, see the Appendix (A).
@@
The optimal trade-off relation in the sequential protocol enables self-testing of crucial quantum components; i) Aparna’s square-shaped pure state preparations on the Bloch sphere, ii) Barun’s measurement instruments defined by Kraus operators with consistent unitary freedom, and iii) Chhanda’s rank-1 projective measurements equivalent to Pauli operators under a unitary transformation. It is to be noted that robust self-testing of states and measurements similar to i) and iii), mentioned above, has been shown earlier both in device-independent [6, 7, 8] and semi device-independent [33, 35] settings. But in such standard settings, certification of quantum instruments is impossible [67, 68]. It is only in the sequential settings where certification of non-projective and unsharp measurements [68, 69] (instruments) [67] can be exhibited. Therefore, in our work we focus on the robust certification of quantum instruments in a SDI way in the next subsection. A thorough robustness of the scheme is shown by bounding the estimation of the sharpness parameter $\eta$, from below and above, considering noise in the state preparation, leading to an improvement over earlier results [67].
@@ III.3 Robust certification of sharpness parameter
In realistic conditions, all quantum systems deviate from the ideal target description, and hence, less than optimal success probabilities are observed, which can’t pinpoint the desired objects. Robustness analysis is the estimation of those objects for the suboptimal observed values. Therefore, any meaningful self-testing argument must come with a robustness analysis. Our above-mentioned self-testing statements in Sec.III.2 are valid only when the optimal trade-off point is achieved, which is not the case in all practical experimental conditions. Here we specifically consider the sharpness parameter of Barun’s measurement, which characterizes his instrument. We show how one can certify the sharpness parameter $\eta$, of Barun’s measurement in the presence of noise, in a more robust way than the previous work [67]. We provide upper and lower bounds on the sharpness parameter for less than optimal observed success probabilities, which merge to the same value in noiseless conditions.
@@
Any binary outcome qubit observable can be expressed in the following way: $M_{y}=g_{y0}\mathbb{I}+\vec{g_{y}}.\vec{\sigma}$. Here, the sharpness parameter of Barun’s measurement is equivalent to the length of the Bloch vector $\vec{g_{y}}$. For simplicity, we take the sharpness parameters of Barun’s two settings to be equal, viz., $\eta\equiv|\vec{g}_{0}|=|\vec{g}_{1}|$. With these specifications, we can now place a lower bound on $\eta$, based on the consideration that the success probability of Barun $P_{AB}^{Q}$, has to be non-classical.
$\eta\geq\sqrt{2}(2P_{AB}^{Q}-1)$ (14)
This bound is non-trivial whenever $P_{AB}^{Q}>1/2$.
On the other hand, from the requirement that $P_{AC}^{Q}$ is non-classical, we can place an upper bound on $\eta$. For obtaining the upper bound of $\eta$, we make use of the optimal choice of the success probability from Eq.(11), given
$P_{AC}^{Q}\leq\frac{1}{2}+\frac{\sqrt{1-\eta^{2}}}{2\sqrt{2}}$ (15)
Inverting the above inequality, leads to,
$\eta\leq\sqrt{8P_{AC}^{Q}-8(P_{AC}^{Q})^{2}-1}$ (16)
@@
From the observed value of $(P_{AB}^{Q},P_{AC}^{Q})$, the pair, one can estimate the sharpness parameter, which lies within the upper and lower bounds. The narrower the gap between the two bounds, the better the certification. Specifically, we consider a noise model where Aparna’s preparation is given by $\rho_{x}^{{}^{\prime}}=p_{1}\rho_{x}+(1-p_{1})\frac{\mathbb{I}}{2}$, and the POVM elements corresponding to the measurement devices of Barun and Chhanda are represented as $M_{y}^{{}^{\prime}}=p_{2}M_{y}+(1-p_{2})\frac{\mathbb{I}}{2}$ and $N_{z}^{{}^{\prime}}=p_{3}N_{z}+(1-p_{3})\frac{\mathbb{I}}{2}$, respectively. Here, $p_{1}$ denotes the visibility of Aparna’s preparation device, $p_{2}$ and $p_{3}$ are the sharpness parameter of Barun’s and Chhanda’s measurement devices, respectively. In Fig. 3, we graphically illustrate the width of the plausible range for the sharpness parameter associated with the underlying noisy measurements.
In a noisy and suboptimal scenario, the sharpness parameter can be estimated within a significantly narrower range using our proposed communication game compared to the estimation based on the sequential QRAC protocol  [67]. For instance, suppose Aparna’s preparations have a visibility of 95%, while Barun’s and Chhanda’s measurements achieve visibilities of 93% and 95%, respectively. Under these conditions, our protocol estimates the sharpness parameter with an error gap of approximately $0.1133$, whereas the method in Ref. [67] leads to a larger error gap of $0.1964$. In Fig. 3(b), for the same visibility values, we plot the gap $(\delta,\delta^{*})$ as a function of $\eta^{\text{Target}}$, where $\eta^{\text{Target}}$ denotes the value of the sharpness parameter that an experimentalist may want to self-test. Detailed calculations and comparison plots for two other noisy scenarios are provided in Fig. 5 of Appendix B.
In Table 1, we present a comparison of the gap between the upper and lower bounds, estimated from the observed figures of merit in our proposed communication game and the protocol without such collaboration [67], for three different sets of noise levels. The gap between the bounds over the entire range of allowed observed success probabilities for the visibility set $(p_{1}=0.95,p_{2}=0.9,p_{3}=0.95)$ for the first set of noise levels is illustrated in Fig. 3(a). The results for the other two noisy scenarios are extracted from Fig. 5 in Appendix B. It can be seen that the protocol for self-testing the sharpness parameter based on our scheme is more robust compared to the one in Ref. [67] for a range of parameter values.